MEASUREMENTS OF LINE POSITIONS AND INTENSITIES OF ^{14}NH_{3} IN THE 1.5 μm REGION

\underline{KEEYOON SUNG}, LINDA R. BROWN, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr.,Pasadena, CA 91109, U.S.A.; XINCHUAN HUANG, SETI Institute, Mountain View, CA 94043, U.S.A.; DAVID W. SCHWENKE, TIMOTHY J. LEE, NASA Ames Research Center, Moffett Field, CA, 94035, U.S.A..

In the atmosphere of outer planets, low mass brown dwarfs, and possibly extrasolar planets, ammonia (NH_{3}) is one of the major opacity sources particularly in the 1.5 μm region (the H-band). However, the spectroscopic information of NH_{3} in the region is completely missing in the HITRAN database. NH_{3} has four infrared active fundamental modes, with the well-known inversion doubling for ν_{2} band, in addition to the usual vibrational degeneracies. Its strong bands, ν_{1}, ν_{3} and 2ν_{4}, dominate the spectrum at 3 μm, while their corresponding overtone and combination bands (e.g., 2ν_{1}, 2ν_{3}, ν_{1}+ν_{3}, ν_{1}+2ν_{4} and ν_{3}+2ν_{4}) are prominent in the 1.5 μm region. As part of an effort to provide a complete set of NH_{3} spectroscopic information in the 1.5 μm region, we are analyzing the laboratory spectra recorded at various temperatures (200 - 299 K) with the McMath-Pierce Fourier transform spectrometer (FTS) on Kitt Peak Observatory in Arizona. Line positions and strengths have been measured from the laboratory spectra, from which lower state energies and quantum assignments are being determined by adopting intensity ratios at two different temperatures and combination differnces. A theoretical IR linelist ^{a} built upon the recent HSL-2 potential energy surface (nonadiabatic corrections included) is complementarily used for the quantum assignments. Preliminary results are presented for ν_{1}+ν_{3}, 2ν_{3}, ν_{1}+2ν_{4} and ν_{3}+2ν_{4} bands and compared with those from early work available.^{b}

aX. Huang, D.W. Schwenke, and T.J. Lee J. Chem. Phys. \underline{134}, (2011) 044320/044321

bThe research described in this paper was performed at the Jet Propulsion Laboratory, California Institute of Technology and the Ames Research Center under contracts with National Aeronautics and Space Administration.